40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 10262     Entry time: Wed Jul 23 11:32:04 2014
Author: Koji 
Type: Update 
Category: LSC 
Subject: IFO warming up 

Alone with the IFO. Started from some conversation with it.

Some ALS trials: Found the Y-end green alignment was terrible. In fact the end green set up is terrible.
Unfixed optics, clipping/fringing in the faraday, unstable suprema mounts which is unnecessarily big.

Eventualy I stopped touching the end alignment. Run ALS to see the stability of the things.
This is a performance confirmation after some touching of the ALS electronics by Manasa/SURFs

The sensing noise levels of the ALSs looks the same as before.

The intensity noise of the transmission was also checked. They are not RIN but very close to RIN
as the DC was the unity for both arms.

The X arm has worse ALS noise level and RIN.
Although I forgot to turn off the HEPA flow at the south bench during the measurement. Gurrr.

Attachment 1: 140722_ALS.pdf  218 kB  | Hide | Hide all
140722_ALS.pdf
ELOG V3.1.3-